Assume the following information:

Current spot rate of British Pound =$1.60

1-year forward rate (as of today) for British Pound =$1.65

Expected spot rate one year from today=$1.67

Rate on 1-year deposits denominated in British Pound =2%

Rate on 1-year deposits denominated in A$ =4%

From the perspective of Australian investors with AUD1, 600 or GBP 1000, what is the rate of return yielded from the covered interest arbitrage??
Answer #1

We can invest GBP 1000 into GBP denominated deposits at interest rate of 2% therefore we have after 1 year GBP denominated deposit equals to 1000(1.02)=GBP 1020

Today we can borrow 1600 AUD at interest rate equals to 4% for a year.

We need to give back AUD 1664 after a year.

We can convert 1020 GBP into AUD using forward rate 1.65 therefore we have 1020*1.65=AUD 1683 after a year

Therefore we can payback AUD 1664 out of AUD 1683 and arbitrage profit equals to 19 AUD

Rate of return is 19/1600=1.1875% per year if investor has no AUD or GBP owned by him at the beginning so if he borrows then his rate of return is 1.1875%

But if he already has AUD 1600 or GBP 1000 then his return is (1683-1600)/1600=83/1600=5.1875%

Rate if return yielded from the covered interest arbitrage is 5.1875%